You are given three investment alternatives to analyze. The cash flows from these three investments are as​ follows:

Investment

End 0f Year

A

B

C

1

$3,000

$1,000

$6,000

2

4,000

1,000

6,000

3

5,000

1,000

(6,000)

4

(6,000)

1,000

(6,000)

5

6,000

5,000

16,000

a.  What is the present value of investment A at an annual discount rate of 15 percent?

b.  What is the present value of investment B at an annual discount rate of 15 percent?

c.  What is the present value of investment C at an annual discount rate of 15 percent?
Answer #1

D8 fc =SUM(D3:07) BC D А 1 A 2 year Cash flows pv@15% Present value 1 $ 3,000.00 0.8696 $ 21 $ 4,000.00 0.7561 $ $ 5,000.00 0

D8 fc =SUM(D3:07) в A 1 A 2 year 31 42 53 64 75 Cash flows 3000 4000 5000 pv@15% =1/1.15 =C3/1.15 =C4/1.15 =C5/1.15 =C6/1.15